At a particular temperature, 12.0 mol of SO_3 is p
Answer #1

convert the given moles in to Molarity

Molarity = moles / volume

intial concentration of SO3 = 12.0 mol / 3L = 4 M

concentration of SO2 at equilibrium = 3.0 mol / 3.0 L = 1 M

construct the ICE table

2SO3 (g) <====> 2SO2 (g) + O2 (g)

I 4 0 0

C -2x +2x +x

E 4-2x +2x +x

from the ICE table equilibrium concewntation of SO2 = 2x

but he has alredy given the equilibrium concentration of SO2 = 1
from these two

2x = 1

x = 1/2 = 0.5

so equilibrium concentrations of

[SO3] = 4-2x = 4 - 2 (0.5) = 4 - 1 = 3

[O2] = x = 0.5 M

now

Kc = [SO2]2 [O2] / [SO3]2

Kc = 32 [ 0.5] / [3]

Kc = 1.5

option A is correct answer